the sum of the 3sides of a triangle is 10(2/3)cm.if two of the sides measure 2(3/4)cm and 4(1/2)cm.find the length of the third side?
Answer by Cromlix(4381)

Hi there,
First add your two sides:
2 (3/4) + 4 (1/2)
= 11/4 + 9/2
= 29/4
Next subtract the answer from the total:
10 (2/3) - 29/4
32/3 - 29/4
= 128 - 87/12
= 41/12
= 3 (5/12) The third side
Hope this helps :-)